# Mind the GAP! The Challenges of Scale in Pixel-based Deep Reinforcement Learning

### v1

@@ Appendix A Experimental Details
Atari Games (Bellemare et al., 2013)
We use the set of 20 games used in Sokar et al. (2025); Obando Ceron* et al. (2024) for direct comparison. The set has the following games: Asterix, SpaceInvaders, Breakout, Pong, Qbert, DemonAttack, Seaquest, WizardOfWor, RoadRunner, BeamRider, Frostbite, CrazyClimber, Assault, Krull, Boxing, Jamesbond, Kangaroo, UpNDown, Gopher, and Hero. This set is used in most of our analysis, nevertheless we provide our main results on the full suite of $60$ games.
Atari100K Games (Łukasz Kaiser et al., 2020)
@@
Appendix B Extra Experiments
B.1 Comparison against max pooling
We compare the performance of global average pooling with global max pooling for Rainbow on the set of 20 games. We find the GAP outperforms max pooling as demonstrated in Figure 12. We hypothesize this is due to GAP’s ability to retain more comprehensive information through its averaging operation.
@@ B.1 Comparison against max pooling
B.2 Performance throughout training

### v2

@@ Appendix B Experimental Details
Atari Games [Bellemare et al., 2013]
We use the set of 20 games used in Sokar et al. [2025], Obando Ceron* et al. [2024] for direct comparison. The set has the following games: Asterix, SpaceInvaders, Breakout, Pong, Qbert, DemonAttack, Seaquest, WizardOfWor, RoadRunner, BeamRider, Frostbite, CrazyClimber, Assault, Krull, Boxing, Jamesbond, Kangaroo, UpNDown, Gopher, and Hero. This set is used in most of our analysis, nevertheless we provide our main results on the full suite of $60$ games.
Atari100K Games [Łukasz Kaiser et al., 2020]
@@
Appendix C Extra Experiments
Figure 12: Sparsification of $\psi$ yields better performance than sparsifying $\phi$ and $\psi$ for 80% sparsity (left) and 95% sparsity (right).
Sparse methods address the bottleneck
Extending our analysis in section 3, we validate our hypothesis on various sparsity levels. Consistent with our main results, sparsification of only $\psi$ yields better performance than sparsifying $\phi$ and $\psi$ across all sparsity levels as shown in Figure 12.
More agents
We evaluate two more algorithms in the sample-efficient regime: DrQ and DrQ($\epsilon$) [Yarats et al., 2021a, Agarwal et al., 2021] on Atari100K. Consistent with our main results, GAP improves performance of scaled networks as shown in Figure 13.
Comparison against max pooling
We compare the performance of global average pooling with global max pooling for Rainbow on the set of 20 games used in our main results. We find the GAP outperforms max pooling as demonstrated in Figure 16. We hypothesize this is due to GAP’s ability to retain more comprehensive information through its averaging operation.
Figure 13: Performance for DrQ (top) and DrQ($\epsilon$) (bottom) [Yarats et al., 2021a, Agarwal et al., 2021] on Atari100K.
Encoder width scaling
We perform additional experiment analyzing the effect of scaling the width of layers in $\phi$ by a factor of 4, while maintaining $\psi$ unscaled. Figure 16 shows that GAP significantly improves performance.
Deeper networks
While our main focus in this work is scaling the width of the network, we also performed some preliminary analysis on scaling the depth of the fully connected layers ($\psi$), exploring architectures with 1, 2, and 3 additional layers. As shown in Figure 16, increasing the depth degrades the performance of the agent. Increasingly, GAP improves performance over the corresponding baseline network of the same size across varying depth.
@@ Appendix C Extra Experiments
Figure 15: Effect of increasing the width of $\phi$. GAP improves performance over the scaled baseline.
Figure 16: Effect of increasing the depth of $\psi$. GAP achieves better performance than the corresponding baseline with the same network size.
Performance throughout training